Find the length of each side of the rectangle so that it has the given perimeter. 2x-5

x
P=50 inches

Answer by checkley71(8403)

P=(2X-5)+2X
50=2X-5+2X
4X=55
X=55/4 SOLUTION
PROOF
50=2*55/4-5+2*55/4
50=110/4-5+110/4
50=220/4-5
50=55-5
50=50